OpenStudy (anonymous):

solve the system by graphing {2x+y=-7 {-3x+2y=14

OpenStudy (anonymous):

y=-2x-7 y=3/2x+7 okay so you would graph that and see where the line crosses which by filling out the graph it would be at x=-4 and y=1

OpenStudy (anonymous):

hmm it says solve by graphing. So im not sure how it is supposed to look . post this in the chat and see if anyone else understands how the answer is supposed to be

OpenStudy (anonymous):

nevermind. I was wrong. Can you show me exactly what you did please

OpenStudy (anonymous):

well I just put the equations in slope intercept form which Im sure you know how to do. And I graphed each equation. And where the 2 lines intersect is where your solution is

OpenStudy (anonymous):

show me every step you did. even the steps of turning it into slope intercept form please

OpenStudy (anonymous):

okay but the you can make your own graph because as you can tell by my draxing I am not a good artist

OpenStudy (anonymous):

2x+y=-7 y=-2x-7 -3x+2y=14 2y=3x+14 y=(3/2)x+7 

OpenStudy (anonymous):

the point where the 2 lines meet in that circle is (-4,1)

OpenStudy (anonymous):

I'm lost. I need every step when it comes to working it before the graphing.. the graphing i get , i just need the work.

OpenStudy (anonymous):

2x+y=-7 y=-2x-7 -3x+2y=14 2y=3x+14 y=(3/2)x+7

OpenStudy (anonymous):

immmm still lost

OpenStudy (anonymous):

hmmm i dunno what i can do differently. Maybe someone else will check this out when more people come online in a few hours and show ya in a better way. Sorry =(

OpenStudy (anonymous):

i just font get how you chnaged the second equation/:

OpenStudy (anonymous):

subtract 3x then divde everything by 2 to get y by its self